# Transition Path Sampling with Boltzmann Generator-based MCMC Moves

### v1

@@ 3.2 Calculating the Path Probability
$\begin{gathered}{\bm{x}}_{i+1}={\bm{x}}_{i}+\Delta t{\bm{v}}_{i+1}\\ {\bm{v}}_{i+1}=\alpha{\bm{v}}_{i}+(1-\alpha)\nabla U({\bm{x}}_{i})+\sqrt{{k}_{B}T(1-\alpha^{2})}{\bm{W}}\text{,}\end{gathered}$ (5)
given the velocity ${\bm{v}}_{i}$, and the molecule’s energy function $U:\mathbb{R}^{d\times 3}\mapsto\mathbb{R}$. $\alpha=\exp(-\Delta t)$ for a time step size $\Delta t$ [^3] Similar to classical fixed length transition path sampling, the timestep size $\Delta t$ is not trivial to choose. We discuss this further in Appendix B. and ${\bm{W}}\sim\mathcal{N}(0,\mathbbm{1})$ corresponds to random motion that is scaled proportional to the Boltzmann constant $k_{B}$ and temperature $T$. Notice that in Langevin dynamics, the only randomness when obtaining ${\bm{x}}_{i+1}$ from ${\bm{x}}_{i}$ given the velocity ${\bm{v}}_{i}$ stems from the Gaussian variable $\sqrt{{k}_{B}T(1-\alpha^{2})}{\bm{W}}$. Thus, the probability density $p({\bm{x}}_{i+1},{\bm{v}}_{i+1}~{}|~{}{\bm{x}}_{i},{\bm{v}}_{i})$ of moving from ${\bm{x}}_{i}$ to ${\bm{x}}_{i+1}$ is that of a Gaussian with mean $\mu={\bm{x}}_{i}+{\Delta t}(\alpha{\bm{v}}_{i}+(1-\alpha)\nabla U({\bm{x}}_{i}))$ and a standard deviation of $\sigma=\Delta t^{2}{k}_{B}T(1-\alpha^{2})$.
@@ 3.3 Latent Space Path Proposal Kernel
Gaussian proposal. From a latent path ${\bm{z}}$, we propose a new path $\tilde{{\bm{z}}}=\{{\bm{z}}_{1}+\mathbf{\epsilon}_{1},\ldots,{\bm{z}}_{l}+\mathbf{\epsilon}_{l}\}$ where $\mathbf{\epsilon}_{1},\ldots,\mathbf{\epsilon}_{l}\sim\mathcal{N}({\bm{0}},\mathbf{\Sigma})$. While this independent noise for each frame makes it unlikely that all frames move coherently and produce high-probability paths, this operation can be performed efficiently and allows for fast exploration of the latent space.
@@ A.1 Latent Path MCMC Algorithm
1 Calculate latent space representation of initial path ${\bm{z}}^{(0)}=\left\{F^{-1}\left({\bm{x}}^{(0)}_{1}\right),\dots,F^{-1}\left({\bm{x}}^{(0)}_{l}\right)\right\}$. for $i\leftarrow 1\ldots N$ do
@@
3             Propose new path in latent space $\tilde{{\bm{z}}}=K\left({\bm{z}}^{(i-1)}\right)$. Compute the proposed path in configuration space $\tilde{{\bm{x}}}=\left\{F\left(\tilde{{\bm{z}}}_{1}\right),\dots,F\left(\tilde{{\bm{z}}}_{l}\right)\right\}$. Compute acceptance probability
@@
Draw a uniformly distributed random number $u\sim\mathcal{U}_{[0,1]}$.
@@ Appendix C Latent Hamiltonian Dynamics Kernel
For this, we perform $L$ leapfrog integration steps with a step size of $\epsilon$. The algorithm initializes the variables
${\bm{m}}^{(0)}_{i}$ $={\bm{m}}_{i}+(\epsilon/2)\nabla p_{Z}({\bm{z}}_{i}),$ (10)
@@
${\bm{z}}^{(t+1)}_{i}$ $={\bm{z}}^{(t)}_{i}+\epsilon{\bm{m}}^{(t)}_{i},$ (11)
${\bm{m}}^{(t+1)}_{i}$ $={\bm{m}}^{(t)}_{i}+\epsilon\nabla p_{Z}({\bm{z}}^{(t+1)}_{i}),$
@@
${\bm{m}}^{*}_{i}$ $={\bm{m}}^{(L)}_{i}-(\epsilon/2)\nabla p_{Z}({\bm{z}}^{(L)}_{i}).$
@@ Appendix D Impact of Variance on Latent Noise Proposal Kernel
The variance of the noise added in latent space is a hyperparameter that will impact the similarity of generated paths. The smaller the variance, the more similar the produced paths will be. In Figure 9, we investigate different noise scales. We can see that a larger variance produces a more diverse ensemble of paths, but the runtime of our approach will increase with it as most proposals will be rejected. This is a trade-off similar to what can be observed in shooting-based TPS, where moves with a vastly different velocity will likely not produce paths with a high probability.
Figure 9: Linear latent proposal kernel with different variances. We compare three different path ensembles sampled by applying noise to the frames in latent space. The noise was sampled from a zero-mean normal distribution with a different $\sigma^{2}$.
P33fBwB4ngcAcByHJpPJl+fn54mD3Gg0NrquXxeLRQAAwzAYj8cwTZPwPH9/sVg8PXweDAauqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v2

@@ 3.2 Calculating the Path Probability
$\begin{gathered}{\bm{x}}_{i+1}={\bm{x}}_{i}+\Delta t{\bm{v}}_{i+1}\\ {\bm{v}}_{i+1}=\alpha{\bm{v}}_{i}-(1-\alpha)\nabla U({\bm{x}}_{i})+\sqrt{{k}_{B}T(1-\alpha^{2})}{\bm{W}}\text{,}\end{gathered}$ (5)
given the velocity ${\bm{v}}_{i}$, and the molecule’s energy function $U:\mathbb{R}^{d\times 3}\mapsto\mathbb{R}$. $\alpha=\exp(-\Delta t)$ for a time step size $\Delta t$ [^3] Similar to classical fixed length transition path sampling, the timestep size $\Delta t$ is not trivial to choose. We discuss this further in Appendix B. and ${\bm{W}}\sim\mathcal{N}(0,\mathbbm{1})$ corresponds to random motion that is scaled proportional to the Boltzmann constant $k_{B}$ and temperature $T$. Notice that in Langevin dynamics, the only randomness when obtaining ${\bm{x}}_{i+1}$ from ${\bm{x}}_{i}$ given the velocity ${\bm{v}}_{i}$ stems from the Gaussian variable $\sqrt{{k}_{B}T(1-\alpha^{2})}{\bm{W}}$. Thus, the probability density $p({\bm{x}}_{i+1},{\bm{v}}_{i+1}~{}|~{}{\bm{x}}_{i},{\bm{v}}_{i})$ of moving from ${\bm{x}}_{i}$ to ${\bm{x}}_{i+1}$ is that of a Gaussian with mean $\mu={\bm{x}}_{i}+{\Delta t}(\alpha{\bm{v}}_{i}-(1-\alpha)\nabla U({\bm{x}}_{i}))$ and a standard deviation of $\sigma=\Delta t^{2}{k}_{B}T(1-\alpha^{2})$.
@@ 3.3 Latent Space Path Proposal Kernel
Gaussian proposal. From a latent path ${\bm{z}}$, we propose a new path $\tilde{{\bm{z}}}=\{{\bm{z}}_{1}+\mathbf{\varepsilon}_{1},\ldots,{\bm{z}}_{l}+\mathbf{\varepsilon}_{l}\}$ where $\mathbf{\varepsilon}_{1},\ldots,\mathbf{\varepsilon}_{l}\sim\mathcal{N}({\bm{0}},\mathbf{\Sigma})$. This procedure is visualized in Figure 3. While this independent noise for each frame makes it unlikely that all frames move coherently and produce high-probability paths, this operation can be performed efficiently and allows for fast exploration of the latent space.
Figure 3: Latent Gaussian noise proposal kernel. We apply independent noise $\varepsilon$ to a latent transition path ${\bm{z}}$. This creates a new proposal path $\tilde{{\bm{z}}}$ which will be accepted or rejected.
@@ A.1 Latent Path MCMC Algorithm
1 Calculate latent space representation of initial path ${\bm{z}}^{(0)}=\left\{F^{-1}\left({\bm{x}}^{(0)}_{1}\right),\dots,F^{-1}\left({\bm{x}}^{(0)}_{l}\right)\right\}$.
2for $i\leftarrow 1\ldots N$ do
@@
4
5            Propose new path in latent space $\tilde{{\bm{z}}}=K\left({\bm{z}}^{(i-1)}\right)$.
6            Compute the proposed path in configuration space $\tilde{{\bm{x}}}=\left\{F\left(\tilde{{\bm{z}}}_{1}\right),\dots,F\left(\tilde{{\bm{z}}}_{l}\right)\right\}$.
7            Compute acceptance probability
@@
8            Draw a uniformly distributed random number $u\sim\mathcal{U}_{[0,1]}$.
@@ Appendix C Latent Hamiltonian Dynamics Kernel
For this, we perform $L$ leapfrog integration steps with a step size of $\varepsilon$. The algorithm initializes the variables
${\bm{m}}^{(0)}_{i}$ $={\bm{m}}_{i}+(\varepsilon/2)\nabla p_{Z}({\bm{z}}_{i}),$ (10)
@@
${\bm{z}}^{(t+1)}_{i}$ $={\bm{z}}^{(t)}_{i}+\varepsilon{\bm{m}}^{(t)}_{i},$ (11)
${\bm{m}}^{(t+1)}_{i}$ $={\bm{m}}^{(t)}_{i}+\varepsilon\nabla p_{Z}({\bm{z}}^{(t+1)}_{i}),$
@@
${\bm{m}}^{*}_{i}$ $={\bm{m}}^{(L)}_{i}-(\varepsilon/2)\nabla p_{Z}({\bm{z}}^{(L)}_{i}).$
This kernel is visualized in Figure 9.
Figure 9: Latent Hamiltonian proposal kernel. We compute the gradient of each point in the latent path ${\bm{z}}$. In combination with a random momentum $m_{2}$, we integrate and determine a new position.
@@ Appendix D Impact of Variance on Latent Noise Proposal Kernel
The variance of the noise added in latent space is a hyperparameter that will impact the similarity of generated paths. The smaller the variance, the more similar the produced paths will be. In Figure 11, we investigate different noise scales. We can see that a larger variance produces a more diverse ensemble of paths, but the runtime of our approach will increase with it as most proposals will be rejected. This is a trade-off similar to what can be observed in shooting-based TPS, where moves with a vastly different velocity will likely not produce paths with a high probability. The likelihood of the paths sampled by different noise levels but the same initial path can be seen in Figure 12. We can observe that a higher noise ratio creates more diverse and more probable paths. However, kernels with a higher variance have a lower acceptance probability.
Figure 11: Gaussian noise latent proposal kernel with different variances. We compare three different path ensembles sampled by applying noise to the frames in latent space. The noise was sampled from a zero-mean normal distribution with a different $\sigma^{2}$.
Figure 12: Gaussian noise latent proposal kernel path likelihood. When comparing different noise scales for the Gaussian noise kernel, we observe that the likelihood of paths scales with the variance.
Appendix E Visualization of Transition
In Figure 13, we visualize a transition of alanine dipeptide.
Figure 13: Transition path. We visualize a transition of alanine dipeptide between the states $C_{5}\leftrightarrow\alpha_{R}$ sampled with the Gaussian noise latent space proposal kernel. We see that the transition looks physical and has no clashes.